OpenStudy (anonymous):

Identify the factors of x2 − 8x − 20.

OpenStudy (theviper):

\[=x^2+2x-10x-20\]\[=x(x+2)-10(x+2)\]\[=(x-10)(x+2)\]

OpenStudy (anonymous):

x^2 - 10x + 2x - 20 (x^2 -10x) + (2x -20) x(x - 10) + 2(x- 10) (x + 2)( x- 10)